find each value. round to the nearest tenth if necessary
1) 80% of 120
120 *80% = 96 or
120 *.80 = 96
2) 115% of 6
6 *115% = 6.9
6 *1.15 = 6.9
3) what percent of 128 is 32??
32/128 = .25 or 25%
4) 3 is what percent of 36?
3/36 = 0.083 or 8.30%
5) 23.7 is 30% of what number?
23.7/.30 = 79
6) 71/2% of what is 12?
71/2% = 35.5%
12/.355 = 33.803
I hope it would help you. Goodluck.

Answer by niccolei(11)

1) 80% of 120
120 *80% = 96 or
120 *.80 = 96
2) 115% of 6
6 *115% = 6.9
6 *1.15 = 6.9
3) what percent of 128 is 32??
32/128 = .25 or 25%
4) 3 is what percent of 36?
3/36 = 0.08 or 8%
5) 23.7 is 30% of what number?
23.7/.30 = 79
6) 71/2% of what is 12?
71/2% = 35.50%
12/.355 = 33.80
I hope it would help you. Goodluck.